# Semidiscrete optimal transport with unknown costs

### v3

@@ 4.1 Properties of the objective
For convenience, let $Z_{j,k}=\left(c_{j}\left(X\right)-g_{j}\right)-\left(c_{k}\left(X\right)-g_{k}\right)$ for any $j,k$. Similarly, let $\bar{\delta}_{j,k}=\delta_{j}-\delta_{k}$. The mapping $z\mapsto\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-z\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>z\right\}}$ is decreasing in $z$. Because $F\left(g+\delta,X\right)\leq c_{j}\left(X\right)-g_{j}-\delta_{j}$ for any $j$ by definition, we then have
$\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-F\left(g+\delta,X\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>F\left(g+\delta,X\right)\right\}}$ (18)
$=$ $\max_{j}\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}$
$=$ $\max_{j}\max_{\ell}\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}\leq c_{\ell}\left(X\right)-g_{\ell},c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}$
$=$ $\max_{j}\max_{\ell}\left(Z_{k,j}-\bar{\delta}_{k,j}\right)\cdot 1_{\left\{Z_{k,i}\leq 0,Z_{k,j}>\bar{\delta}_{k,j}\right\}},$
$\geq$ $\max_{j}\left(Z_{k,j}-\bar{\delta}_{k,j}\right)\cdot 1_{\left\{\bar{\delta}_{k,j}<Z_{k,j}\leq 0\right\}},$
where (18) follows by taking $\ell=j$ in the inner maximum. Therefore,
$\mathbb{E}\left(T_{2}\right)$$\geq$ $\sum^{K}_{k=1}\mathbb{E}\left(\max_{j}\left(Z_{k,j}-\bar{\delta}_{k,j}\right)\cdot 1_{\left\{\bar{\delta}_{k,j}<Z_{k,j}\leq 0\right\}}\right)$
$\geq$ $\sum^{K}_{k=1}\max_{j}\mathbb{E}\left(\left(Z_{k,j}-\bar{\delta}_{k,j}\right)\cdot 1_{\left\{\bar{\delta}_{k,j}<Z_{k,j}\leq 0\right\}}\right).$
We now consider two cases. In the first case, $\bar{\delta}_{k,j}<-\kappa_{1}$ and we can use Assumption 3.2(i) to obtain
$\mathbb{E}\left(\left(Z_{k,j}-\bar{\delta}_{k,j}\right)\cdot 1_{\left\{\bar{\delta}_{k,j}<Z_{k,j}\leq 0\right\}}\right)$ $=$ $\int^{0}_{\bar{\delta}_{k,j}}\left(z-\bar{\delta}_{k,j}\right)P\left(Z_{k,j}\in dz\right)$
$\geq$ $\int^{0}_{-\kappa_{1}}\left(z-\bar{\delta}_{k,j}\right)P\left(Z_{k,j}\in dz\right)$
[…]
$=$ $\frac{1}{2}\kappa_{1}\kappa_{2}\left(-2\bar{\delta}_{k,j}-\kappa_{1}\right)$
$>$ $\frac{1}{2}\kappa^{2}_{1}\kappa_{2}.$
In the second case, $-\kappa_{1}\leq\bar{\delta}_{k,j}\leq 0$ and
[…]
$=$ $\frac{1}{2}\kappa_{2}\bar{\delta}^{2}_{k,j}.$
Combining the two cases yields
@@
$\mathbb{E}\left(T_{2}\right)$$\geq$ $\frac{\kappa_{2}}{2}\sum^{K}_{k=1}\max_{j}\min\left\{\kappa^{2}_{1},\bar{\delta}^{2}_{k,j}\right\}1_{\left\{\bar{\delta}_{j,k}\geq 0\right\}}$ (19)
$\geq$ $\frac{\kappa_{2}}{2K}\sum^{K}_{k=1}\sum^{K}_{j=1}\min\left\{\kappa^{2}_{1},\bar{\delta}^{2}_{k,j}\right\}1_{\left\{\bar{\delta}_{j,k}\geq 0\right\}}.$
@@
$\mathbb{E}\left(T_{2}\right)\geq\frac{\kappa_{2}}{2K}\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\},$
@@
$\mathbb{E}\left(T_{1}+T_{2}\right)\geq\delta^{\top}p+\frac{1}{2K}\kappa_{2}\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\},$
@@
$\delta^{\top}\left(p^{g^{*}+\delta}-p\right)\geq\frac{\kappa_{2}}{2K}\min\left\{\|\delta\|^{2}_{2},\kappa^{2}_{1}\right\}.$ (20)
@@
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\kappa^{2}_{1}\right\}.$
To see this, consider two cases. First, suppose that there is some $j$ for which $\left|\delta_{j}\right|\geq\kappa_{1}$. Then,
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\}\geq\min\left\{\delta^{2}_{j},\kappa^{2}_{1}\right\}=\kappa^{2}_{1}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\kappa^{2}_{1}\right\}.$
Second, suppose that $\left|\delta_{k}\right|<\kappa_{1}$ for all $k$. Then,
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\}=\sum^{K}_{k=1}\delta^{2}_{k}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\kappa^{2}_{1}\right\}.$
@@
Let Assumptions 3.1-3.2 hold. There exists a constant $C>0$ such that $P\left(\|\delta^{n}\|_{2}>2\right)\leq\frac{C}{n^{9}}$.
@@
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha}{n+1}\left(\zeta^{n+1}-p^{n}\right)^{\top}\delta^{n}-\frac{2\alpha C_{1}}{n+1}\min\left\{\|\delta^{n}\|^{2}_{2},\kappa^{2}_{1}\right\},$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)\leq\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\kappa^{2}_{1}\right\}.$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\|\delta^{n}\|^{2}_{2}\min\left\{1,\frac{\kappa^{2}_{1}}{\|\delta^{n}\|^{2}}\right\}$
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\|\delta^{n}\|^{2}_{2}\min\left\{1,\frac{\kappa^{2}_{1}}{C^{2}_{0}}\right\}$
$=$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}.$
@@
$\leq$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}1_{\left\{\|\delta^{n}\|_{2}>C_{0}\right\}}\mid\mathcal{F}^{n}\right)$
$\leq$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\alpha C_{2}\log\left(n+1\right)1_{\left\{\|\delta^{n}\|_{2}>C_{0}\right\}},$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\right)\leq\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{C^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\alpha C_{2}\log\left(n+1\right)\frac{C}{n^{9}}.$
@@
5 Analysis for unknown costs
As in Section 3.2, we will focus on the case of linear costs $c_{k}\left(X\right)=\beta^{\top}_{k}X$ where (15) is used for estimation. The Appendix presents a fully rigorous extension to the case of more general link functions $\chi$. The analysis in that extension will be very similar, with some additional technical complications in several steps.
We proceed similarly to Section 4. Section 5.1 proves analogous properties to those in Section 4.1 for the case of unknown costs. Section 5.2 outlines the proof of the main rate results, with some intermediate steps moved to the Appendix. The core of the proof is the derivation of new concentration inequalities for both the SA error and the estimation error. The rates of these inequalities allow us to place a sub-exponential gap between periods of forced exploration.

### v4

@@ 4.1 Properties of the objective
The mapping $z\mapsto\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-z\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>z\right\}}$ is decreasing in $z$. Because $F\left(g+\delta,X\right)\leq c_{j}\left(X\right)-g_{j}-\delta_{j}$ for any $j$ by definition, we then have
$\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-F\left(g+\delta,X\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>F\left(g+\delta,X\right)\right\}}$
$\geq$ $\frac{1}{K}\sum^{K}_{j=1}\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}.$
Consequently,
$T_{2}\geq\frac{1}{K}\sum^{K}_{k=1}\sum^{K}_{j=1}\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}.$ (18)
We now seek to bound the expectation of each term in (18). For convenience, let $Z_{j,k}=\left(c_{j}\left(X\right)-g_{j}\right)-\left(c_{k}\left(X\right)-g_{k}\right)$ for any $j,k$. Similarly, let $\bar{\delta}_{j,k}=\delta_{j}-\delta_{k}$. We write
$\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}$
$=$ $\left(\bar{\delta}_{j,k}-Z_{j,k}\right)\cdot 1_{\left\{\min_{\ell\neq j}Z_{\ell,k}\geq 0,\;0\leq Z_{j,k}<\bar{\delta}_{j,k}\right\}}.$
Taking expectations, we obtain
$\mathbb{E}\left\{\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}\right\}$
$=$ $\mathbb{E}\left\{\left(\bar{\delta}_{j,k}-Z_{j,k}\right)\cdot 1_{\left\{0\leq Z_{j,k}<\bar{\delta}_{j,k}\right\}}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq 0\mid Z_{j,k}\right)\right\}$
$=$ $\int^{\bar{\delta}_{j,k}}_{0}\left(\bar{\delta}_{j,k}-z\right)P\left(\min_{\ell\neq j}Z_{\ell,k}\geq 0\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)$
$\geq$ $\frac{1}{2}\bar{\delta}_{j,k}\int^{\frac{\bar{\delta}_{j,k}}{2}}_{0}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq 0\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)$ (20)
$\geq$ $\frac{1}{2}\bar{\delta}_{j,k}\int^{\frac{\bar{\delta}_{j,k}}{2}}_{0}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)$ (21)
where (4.1) follows by the tower property, and (20) follows because $\bar{\delta}_{j,k}-z\geq 0$ for $z\in\left[0,\bar{\delta}_{j,k}\right]$.
Using Lemma 3.1(ii), and taking $0<\kappa<\kappa^{\prime}_{1}$, we write
$\kappa^{\prime}_{4}$ $\leq$ $P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid 0\leq Z_{j,k}\leq\kappa\right)$ (22)
$=$ $\frac{\int^{\kappa}_{0}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)}{\int^{\kappa}_{0}P\left(Z_{j,k}\in dz\right)},$
and, since $P\left(Z_{j,k}\in dz\right)\geq\kappa^{\prime}_{2}$ on $0\leq z\leq\kappa^{\prime}_{1}$, (22) yields
$\int^{\kappa}_{0}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)\geq\kappa\cdot\kappa^{\prime}_{2}\kappa^{\prime}_{4}.$ (23)
Then, returning to (20), we consider two cases. First, suppose that $0<\bar{\delta}_{j,k}<\kappa^{\prime}_{1}$. Then, (23) can be applied directly, yielding
$\mathbb{E}\left\{\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}\right\}$ (24)
$\geq$ $\frac{1}{4}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\bar{\delta}^{2}_{j,k}.$
In the second case, we consider $\bar{\delta}_{j,k}\geq\kappa^{\prime}_{1}$. The mapping
$x\mapsto\left(x-Z_{j,k}\right)\cdot 1_{\left\{0\leq Z_{j,k}<x\right\}}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq 0\mid Z_{j,k}\right)$
is increasing on the interval $\left[\kappa^{\prime}_{1},\infty\right)$. Therefore, repeating the arguments used to derive (21), and applying (23) again, we obtain
$\mathbb{E}\left\{\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}\right\}$ (25)
$\geq$ $\frac{1}{2}\kappa^{\prime}_{1}\int^{\frac{\kappa^{\prime}_{1}}{2}}_{0}P\left(\min_{\ell\neq j}Z_{\ell,k}\geq\kappa^{\prime}_{1}\mid Z_{j,k}=z\right)P\left(Z_{j,k}\in dz\right)$
$\geq$ $\frac{1}{4}\left(\kappa^{\prime}_{1}\right)^{2}\kappa^{\prime}_{2}\kappa^{\prime}_{4}.$
Combining (24) and (25), we have
$\mathbb{E}\left\{\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}\right\}$ (26)
$\geq$ $\frac{1}{4}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\min\left\{\bar{\delta}^{2}_{j,k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}$
for any $\bar{\delta}_{j,k}>0$. Since the left-hand side of (26) is always positive, we have
$\mathbb{E}\left\{\left[\left(c_{k}\left(X\right)-g_{k}-\delta_{k}\right)-\left(c_{j}\left(X\right)-g_{j}-\delta_{j}\right)\right]\cdot 1_{\left\{c_{k}\left(X\right)-g_{k}=F\left(g,X\right),c_{k}\left(X\right)-g_{k}-\delta_{k}>c_{j}\left(X\right)-g_{j}-\delta_{j}\right\}}\right\}$ (27)
$\geq$ $\frac{1}{4}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\min\left\{\bar{\delta}^{2}_{j,k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}\cdot 1_{\left\{\bar{\delta}_{j,k}\geq 0\right\}}$
for any value (positive or negative) of $\bar{\delta}_{j,k}$.
We may now derive the desired result. Combining (18) and (27), we have
@@
$\mathbb{E}\left(T_{2}\right)\geq\frac{1}{4K}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\sum^{K}_{k=1}\sum^{K}_{j=1}\min\left\{\bar{\delta}^{2}_{j,k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}\cdot 1_{\left\{\bar{\delta}_{j,k}\geq 0\right\}}.$ (28)
@@
$\mathbb{E}\left(T_{2}\right)\geq\frac{1}{4K}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\},$
@@
$\mathbb{E}\left(T_{1}+T_{2}\right)\geq\delta^{\top}p+\frac{1}{4K}\kappa^{\prime}_{2}\kappa^{\prime}_{4}\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\},$
@@
$\delta^{\top}\left(p^{g^{*}+\delta}-p\right)\geq\frac{\kappa^{\prime}_{2}\kappa^{\prime}_{4}}{4K}\min\left\{\|\delta\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\}.$ (29)
@@
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}.$
To see this, consider two cases. First, suppose that there is some $j$ for which $\left|\delta_{j}\right|\geq\kappa^{\prime}_{1}$. Then,
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}\geq\min\left\{\delta^{2}_{j},\left(\kappa^{\prime}_{1}\right)^{2}\right\}=\left(\kappa^{\prime}_{1}\right)^{2}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}.$
Second, suppose that $\left|\delta_{k}\right|<\kappa^{\prime}_{1}$ for all $k$. Then,
$\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}=\sum^{K}_{k=1}\delta^{2}_{k}\geq\min\left\{\sum^{K}_{k=1}\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}.$
@@
Let Assumption 3.1 hold. There exists a constant $C>0$ such that $P\left(\|\delta^{n}\|_{2}>2\right)\leq\frac{C}{n^{9}}$.
@@
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha}{n+1}\left(\zeta^{n+1}-p^{n}\right)^{\top}\delta^{n}-\frac{2\alpha C_{1}}{n+1}\min\left\{\|\delta^{n}\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\},$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)\leq\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\}.$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$ $\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\|\delta^{n}\|^{2}_{2}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{\|\delta^{n}\|^{2}}\right\}$
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha C_{1}}{n+1}\|\delta^{n}\|^{2}_{2}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{C^{2}_{0}}\right\}$
$=$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}.$
@@
$\leq$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}1_{\left\{\|\delta^{n}\|_{2}>C_{0}\right\}}\mid\mathcal{F}^{n}\right)$
$\leq$ $\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{C^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\alpha C_{2}\log\left(n+1\right)1_{\left\{\|\delta^{n}\|_{2}>C_{0}\right\}},$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\right)\leq\left(1-\frac{2\alpha C_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{C^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}+\alpha C_{2}\log\left(n+1\right)\frac{C}{n^{9}}.$
@@
5 Semi-myopic learning: technical analysis
Finally, we may prove our main result (Theorems 3.2-3.3). As in Section 3.2, we will focus on the case of linear costs $c_{k}\left(X\right)=\beta^{\top}_{k}X$ where (15) is used for estimation. The Appendix presents a fully rigorous extension to the case of more general link functions $\chi$. The analysis in that extension will be very similar, with some additional technical complications in several steps.
The order of the arguments is similar to Section 4. Section 5.1 proves analogous properties to those in Section 4.1 for the case of unknown costs. Section 5.2 outlines the proof of the main rate results, with some intermediate steps moved to the Appendix. The core of the proof is the derivation of new concentration inequalities for both the SA error and the estimation error. The rates of these inequalities allow us to place a sub-exponential gap between periods of forced exploration.